# Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions

### v1

@@ 3.1 Global Task Token Learning
Firstly, we flatten the shared feature $\boldsymbol{F_{s}}$ into feature tokens with shape $\mathbb{R}^{C\times(hw)}$, and use each global task token to query the shared feature to obtain each task feature $\boldsymbol{F_{i}}$ accordingly. Then, to conduct inter-task learning, all global task tokens are used to produce cross-task affinities that explicitly guide the learning process. The cross-task affinity map $\boldsymbol{A}\in\mathbb{R}^{T\times T}$ is calculated as:
$\boldsymbol{A}=\operatorname{Softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top}),$ (1)
@@
where $\boldsymbol{Q}$ and $\boldsymbol{K}$ are individual linear projection of concatenated global tokens $\boldsymbol{\Theta}=[\boldsymbol{\theta_{1}};\boldsymbol{\theta_{2}};\cdots;\boldsymbol{\theta_{T}}]^{\top}$, in which []delimited-[][$\cdot$][ ⋅ ] indicates the concatenation. After affinity matrix $\boldsymbol{A}$ is calculated, it is used to conduct affine combinations of task features and global task tokens respectively:
@@
Afterward, the updated tokens and features with cross-task information are involved in the intra-task learning, where we first concatenate every corresponding task token and feature, and perform self-attention on the spatial dimension among each token-feature pair $\left[\boldsymbol{\theta_{i}^{\prime}};\boldsymbol{F_{i}^{\prime}}\right]\in\mathbb{R}^{C\times(hw+1)}$. The global task tokens will further learn more specific and discriminative task representations during this process, and representative task tokens will in turn enhance the feature quality as well. Followingly, we discover pseudo feature supervision with the aid of well-learned global task tokens $\boldsymbol{\theta_{i}^{\prime}}$, and this process will be discussed in Sec. 3.3.
@@
Additionally, for multi-scale backbone features, directly fusing them ignores the various granularity of task representations maintained at different scales. Thus, for multi-scale image backbone, we further propose Multi-scale Global Task Token Learning in order to learn comprehensive multi-scale task relations. The proposed method involves inter-task learning separately at each scale, and then the multi-scale features and tokens are fused before intra-task learning. In this way, the global task tokens gain richer cross-task relations at different scales and are able to maintain stronger representations. We will illustrate this part in detail in the supplementary material.
3.2 Fine-grained Task Token Learning
After the global task tokens are learned, we further propose to conduct feature-token interaction at a finer spatial granularity, which takes advantage of various representations in global task tokens and boosts task label discovery process. The process is illustrated in Fig. 4.
Firstly, we jointly project each updated task token $\boldsymbol{\theta_{i}^{\prime}}\in\mathbb{R}^{C}$ and feature $\boldsymbol{F_{i}^{\prime}}\in\mathbb{R}^{C\times h\times w}$ into the prediction space with finer granularity. For features, this can be easily achieved by applying a linear convolution layer, and we denote the fine-grained task features as $\boldsymbol{G_{i}}\in\mathbb{R}^{C_{p}\times H\times W}$, where $C_{p}$ indicates the prediction dimension. For tokens, we denote the projected fine-grained tokens as $\boldsymbol{\varphi_{i}}\in\mathbb{R}^{C_{p}\times C}$, we hope every $1\times C$ vector inside it can represent one category distribution over the spatial dimension. The simplest way is to project each $\boldsymbol{\theta_{i}^{\prime}}$ with a Multi-Layer Perceptron (MLP), which can be described as: $\boldsymbol{\varphi_{i}}=\operatorname{MLP_{i}}(\boldsymbol{\theta_{i}})^{\top},i=1,2,\cdots,T$. However, since there are no direct supervision imposed during this process to distinguish the learned fine-grain tokens, the MLPs will tend to degenerate and perform linearly correlated outputs, which prevents the fine-grained tokens from learning discriminative task-specific representations. To alleviate this problem, we propose to use Orthogonal Embeddings (OE) to serve as priors and aid the learning process of fine-grained tokens.
Figure 4: Illustration of Token &\&& Feature Projection and the Fine-grained Task Token Learning. Different tasks have the same structure, and we take one task as example. We firstly project the fine-grained feature $\boldsymbol{G_{i}}$ from the updated task feature $\boldsymbol{F_{i}^{\prime}}$, and derive fine-grained task tokens $\boldsymbol{\varphi_{i}}$ from the updated global task tokens $\boldsymbol{\theta_{i}^{\prime}}$. Then, we produce final task predictions with pseudo label discovery by conducting interactions between $\boldsymbol{G_{i}}$ and $\boldsymbol{\varphi_{i}}$.
In order to learn discriminative task representations, the vectors in the fine-grained token should be far from each other to represent meaningful and distinguishable task category information, so we use a group of orthogonal basis in $\mathbb{R}^{C_{p}}$ to serve as the embedding for MLP input, denoting as $\boldsymbol{o}\in\mathbb{R}^{C_{p}\times C_{p}}$. These OE are projected into the feature space by linear projection matrix $\boldsymbol{W_{i}^{o}}\in\mathbb{R}^{C\times C_{p}}$, and then added with the global token before being fed into the MLP. The process can be described as:
$\boldsymbol{\varphi_{i}}=\operatorname{MLP_{i}}(\boldsymbol{W_{i}^{o}}\times\boldsymbol{o}+[\underbrace{\boldsymbol{\theta_{i}^{\prime}};\boldsymbol{\theta_{i}^{\prime}};\cdots;\boldsymbol{\theta_{i}^{\prime}}}_{C_{p}}])^{\top}.$ (3)
With the prior information from $\boldsymbol{o}$, the MLPs are capable of keeping the distance between the vectors in $\boldsymbol{\varphi_{i}}$ far from each other in the feature space, which makes it able to learn information that distinguishes between task categories as well as inherit the global representations in $\boldsymbol{\theta_{i}^{\prime}}$.
Subsequently, we exploit the fine-grained task tokens $\boldsymbol{\varphi_{i}}$ for the fine-grained token learning process. Normally, $\boldsymbol{G_{i}}$ is noisy and low-confident on unlabeled data due to the lack of supervision, which leads to inaccurate predictions. Since we are treating all of the tasks as pixel-wise classification, we propose to use fine-grained task tokens to encourage $\boldsymbol{G_{i}}$ to produce high-confident logits, and enhance the quality of pseudo labels:
$\boldsymbol{G_{i}^{\prime}}=\operatorname{Conv_{i}^{3\times 3}}\left(\boldsymbol{G_{i}}\odot\operatorname{Softplus}\left(\boldsymbol{\varphi_{i}}\times\boldsymbol{G_{i}}\right)\right),$ (4)
where $\operatorname{Softplus}(x)=\log(1+\exp(x))$, and $\operatorname{Softplus}(\boldsymbol{\varphi_{i}}\times\boldsymbol{G_{i}})$ gives pixel-wise positive score masks on each channel of $\boldsymbol{G_{i}}$. Since $\boldsymbol{\varphi_{i}}$ inherits global task representations from $\boldsymbol{\theta_{i}^{\prime}}$, which will perform more robustly on unlabeled data, and aid the production of distinguished logits score in the updated feature $\boldsymbol{G_{i}^{\prime}}$, as well as high-confident final task predictions. Similar to the previous stage, we also conduct pseudo label discovery after the fine-grained tokens are learned, which will be discussed in detail in the next section.
3.3 Hierarchical Label Discovery and Multi-task Optimization
$\mathcal{L}_{s}=\frac{1}{T}\sum_{i=1}^{T}\left(\alpha_{i}\operatorname{L_{i}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right)\right),$ (5)
We firstly train the multi-task model with HiTTs jointly with $\operatorname{L_{i}}$ only to achieve convergence on labeled data, then we utilize both hierarchies of tokens to discover feature-level and prediction-level pseudo supervision. As we mentioned in Sec. 3.1, before the updated tokens and features are fed into the next stage, we conduct the feature supervision discovery to excavate feature-level supervision signals for unlabeled tasks. As shown in Fig. 3, we use the updated global task tokens to query each pixel feature on every task and produce confidence mask $\boldsymbol{M_{i}^{f}}=\operatorname{Sigmoid}(\boldsymbol{\theta_{i}^{\prime}}^{\top}\times\boldsymbol{F_{i}^{\prime}})$. Since $\boldsymbol{\theta_{i}}$ is globally learned on all task features, in $\boldsymbol{M_{i}^{f}}$, higher scores indicate that the pixel features have a higher response to task $i$, which should be further used to prove task supervision. Thus we use $\boldsymbol{M_{i}^{f}}$ to serve as a soft confidence mask for pixel-wise feature supervision loss:
$\mathcal{L}_{f}$ $=\frac{1}{T}\sum_{i=1}^{T}\left(\alpha_{i}\operatorname{L_{mse}}\left(\boldsymbol{F_{i}^{\prime}},\boldsymbol{\tilde{F_{i}}}\right)\right.$
$\left.+(1-\alpha_{i})\boldsymbol{M_{i}^{f}}\odot\operatorname{L_{mse}}\left(\boldsymbol{F_{i}^{\prime}},\boldsymbol{\tilde{F_{i}}}\right)\right),$ (6)
where $\tilde{\boldsymbol{F_{i}}}$ represents the offline saved features which serve as pseudo supervision signals for unsupervised task features. The $\odot$ represents element-wise multiplication, $\operatorname{L_{mse}}$ is the mean squared error loss for feature distance measurement. For the feature loss on labeled task (first item in Eq 3.3), we regard all pixel features from $\boldsymbol{\tilde{F_{i}}}$ as valid since they are supervised by ground-truth label, while for unlabeled task (second item in Eq 3.3), we use $\boldsymbol{M_{i}^{f}}$ encourage high-confidence pixel features and depress low-confidence ones.
Afterward, we also conduct pseudo label discovery with the aid of well-learned fine-grained task tokens $\boldsymbol{\varphi_{i}}$ as mentioned in Sec. 3.2. We directly produce pseudo labels from $\boldsymbol{G_{i}^{\prime}}$: $\boldsymbol{\tilde{Y_{i}}}=\operatorname{Argmax}\left(\operatorname{Softmax}\left(\boldsymbol{G_{i}^{\prime}}\right)\right)$, along with binary masks to select high confidence pixel pseudo labels: $\boldsymbol{M_{i}^{p}}=\operatorname{Max}\left(\operatorname{Softmax}\left(\boldsymbol{G_{i}^{\prime}}\right)\right)>\tau_{i}$, where $\tau_{i}$ is a threshold used to produce binary masks. The loss for pseudo label supervision can be written as:
$\mathcal{L}_{p}=\frac{1}{T}\sum_{i=1}^{T}\left((1-\alpha_{i})\boldsymbol{M_{i}^{p}}\odot\operatorname{L_{i}}\left(\boldsymbol{\hat{Y_{i}}},\boldsymbol{\tilde{Y_{i}}}\right)\right),$ (7)
Finally, we sum all of the losses in Eq 5,  3.3 and  7 to supervise all task features and predictions. The overall losses to optimize the model can be described as:
$\mathcal{L}$ $=\mathcal{L}_{s}+\mathcal{L}_{p}+\mathcal{L}_{f}$ (8)
$=\frac{1}{T}\sum_{i=1}^{T}\left[\alpha_{i}\left(\operatorname{L_{i}}(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}})+\operatorname{L_{mse}}(\boldsymbol{F_{i}^{\prime}},\boldsymbol{\tilde{F_{i}}})\right)\right.$
$+\left.(1-\alpha_{i})\left(\boldsymbol{M_{i}^{p}}\odot\operatorname{L_{i}}(\boldsymbol{\hat{Y_{i}}},\boldsymbol{\tilde{Y_{i}}})+\boldsymbol{M_{i}^{f}}\odot\operatorname{L_{mse}}(\boldsymbol{F_{i}^{\prime}},\boldsymbol{\tilde{F_{i}}})\right)\right]$
where the second row represents all prediction-level supervision and the third row represents all feature-level supervision.
4 Experiment
4.1 Experimental Setup
NYUD-v2. NYUD-v2 silberman2012indoor contains 795 and 654 RGB-D indoor scene images for training and testing respectively. We use the 13-class semantic annotations which is defined in couprie2013indoor , the truth depth annotations recorded by Microsoft Kinect depth camera, and surface normal annotations which are produced in eigen2015predicting . Following the setting in li2022learning ; liu2019end , we use $288\times 384$ image resolution to speed up training. We use Adam optimizer with a learning rate of $1\times 10^{-4}$, and train all models for 400 epochs with batch size $8$. We update the learning rate every $100$ epoch with $\gamma=0.5$ as the multiplying factor.
Cityscapes. Cityscapes cordts2016cityscapes contains 2975 and 500 street-view images for training and testing respectively. We used the projected 7-class semantic annotations from liu2019end , and disparity maps to serve as depth annotations. Following the setting in li2022learning ; liu2019end , we use $128\times 256$ image resolution to speed up training. The optimizer and learning rate scheduler are set as the same as NYUD-v2.
@@ 4.1 Experimental Setup
PASCAL-Context. PASCAL-Context everingham2009pascal contains 4998 and 5105 images for training and testing respectively, which also have pixel-level annotations for semantic segmentation, human-parts segmentation and semantic edge detection. Additionally, we also consider surface normal estimation and saliency detection distilled by maninis2019attentive . We use Adam optimizer with learning rate $1\times 10^{-4}$, and weight decay $1\times 10^{-4}$, and train for $20000$ steps with batch size $6$. We update the learning rate by polynomial scheduler and $\gamma=0.9$ for the power factor.
Model Setting. Following li2022learning , we use SegNet badrinarayanan2017segnet for NYUD-v2 and Cityscapes, ResNet-18 he2016deep for PASCAL-Context as the backbone of our single task learning (STL) baselines, and the multi-task baseline (MTL) is built from it, which consists of a shared backbone encoder and several task-specific decoding heads. For the learning of HiTTs, we follow bruggemann2021exploring ; li2018monocular ; zeisl2014discriminatively , and perform a discrete quantization of the label space of continuous regression tasks such as Depth. and Normal. This discrete quantization does not contribute to multi-task learning performance as analysed in bruggemann2021exploring , so we ensure a fair comparison with other works. The network with HiTTs is first jointly trained only on labeled data to obtain high-quality task representations and produce pseudo labels, and then follow the self-training procedure in xie2020self , we optimize the model with labels or pre-produced pseudo labels on all tasks and train from scratch, and the feature-level supervision is also applied during the training period.
Data Preparation. We follow the setting of li2022learning ; liu2019end to process training data, and form two partially annotated settings li2022learning : (i) one-label: for each input image, it is only associated with one task annotation; (ii) random-labels: each image has at least one and at most $N-1$ tasks with corresponding task annotations, in the set of $N$ tasks. Additionally, we provide two extra settings to further validate the effectiveness of our method: (iii) full-labels: each image has labels on every task, which is the usual multi-task learning setting; (iv) few-shot: one certain task has only very few labels while the other tasks are fully supervised.
Training Pipeline. We first train the multi-task model with HiTTs on all labeled task data. Then the weights of the network and tokens are fixed, and used to produce hierarchical supervision on both feature and prediction spaces. We produce the pseudo label in an offline manner according to xie2020self , which is labeling on clean image without data augmentation, and training on augmented images and pseudo labels to enforce consistent predictions. In xie2020self , this method is only applied to classification tasks while we extend the utilization to general dense prediction tasks. After the pseudo labels are produced, we use them along with the ground-truth labels to jointly train the multi-task model from scratch. Additionally, we also use the pseudo feature supervision produced by this pre-trained multi-task model for feature regularization during the optimization process. Both hierarchies of the discovered supervision signals ensure that all task predictions will obtain pixel-wise supervision for multi-task optimization to gain better generalization ability on unlabeled data.
Evaluation Metrics. We have briefly introduced our evaluation metrics for multiple dense prediction tasks in Sec.4.1. We provide a more detailed description as follows: (i) mIoU: mean intersection over union; (ii) pAcc: per-pixel accuracy; (iii) AbS or AbR: absolute error or absolute-relative error; (iv) rmse: root mean square error (for Normal. we calculate the mean square error of the predicted angles with the ground-truths); (v) mErr: mean of angle error; (vi) odsF: optimal dataset F-measure martin2004learning ; (vii) threshold: for surface normal estimation, we calculate the proportion of pixels with angle error smaller than three thresholds $\eta\in\left\{11.25^{\circ},22.50^{\circ},30^{\circ}\right\}$. Additionally, to better evaluate the proposed method, we also consider using $\Delta_{MTL}$ proposed by vandenhende2021multi to evaluate the overall improvement of the multi-task performances of all the tasks, which is defined as:
$\Delta_{MTL}=\sum_{N}(-1)^{l_{t}}\left(M_{m,t}-M_{s,t}\right)/M_{s,t}$ (9)
Figure 5: Comparisons with SoTA works on NYUDv2. Ours shows both clear semantic boundaries and accurate geometry estimations, indicating the effectiveness of cross-task feature-token interactions.

### v2

@@ A.2. Global Task Token Learning
Firstly, we flatten the shared feature $\boldsymbol{F_{s}}$ into feature tokens with shape $\mathbb{R}^{C\times(hw)}$, and use each global task token to query the shared feature to obtain each task feature $\boldsymbol{F_{i}}$ accordingly. Then, to conduct inter-task attention, all global task tokens are used to produce cross-task affinities that explicitly guide the learning process. The cross-task affinity map $\boldsymbol{A}\in\mathbb{R}^{T\times T}$ is calculated as:
(5) $\boldsymbol{A}=\operatorname{Softmax}(\boldsymbol{Q}\times\boldsymbol{K}^{\top}),$
@@
where $\boldsymbol{Q}$ and $\boldsymbol{K}$ are individual linear projection of concatenated global tokens $\boldsymbol{\Theta}=[\boldsymbol{\theta_{1}};\boldsymbol{\theta_{2}};\cdots;\boldsymbol{\theta_{T}}]^{\top}$, in which [][$\cdot$][ ⋅ ] indicates the concatenation. After affinity matrix $\boldsymbol{A}$ is calculated, it is used to conduct affine combinations of task features and global task tokens respectively:
@@
Afterward, the updated tokens and features with cross-task information are involved in the intra-task attention, where we first concatenate every corresponding task token and feature, and perform self-attention on the spatial dimension among each token-feature pair $\left[\boldsymbol{\theta_{i}^{\prime}};\boldsymbol{F_{i}^{\prime}}\right]\in\mathbb{R}^{C\times(hw+1)}$. The global task tokens will further learn more specific and discriminative task representations during this process, and representative task tokens will in turn enhance the feature quality as well. Followingly, we discover pseudo feature supervision with the aid of well-learned global task tokens $\boldsymbol{\theta_{i}^{\prime}}$.
@@
Additionally, for multi-scale backbone features, directly fusing them ignores the various granularity of task representations maintained at different scales. Thus, for multi-scale image backbone, we further propose Multi-scale Global Task Token Learning in order to learn comprehensive multi-scale task relations. The proposed method involves inter-task attention separately at each scale, and then the multi-scale features and tokens are fused before intra-task attention. In this way, the global task tokens gain richer cross-task relations at different scales and are able to maintain stronger representations. We will illustrate this part in detail in Sec. A.4.
Figure 8. Illustration of detail designs of: (i) Inter-Task Attention: predicting cross-task token affinities $\boldsymbol{A}$ from the global task tokens. (ii) Intra-Task Attention: conducting self-attention between task features and tokens.
A.3. Discrete Quantization and Task Losses
In Sec. 3 of the body part, we have discussed how to learn HiTTs. For continuous regression tasks, such as Depth and Normal, we first need to perform a discrete quantization of the label space to provide discriminative supervision for tokens. The goal for quantization is to assign meaningful category bins to each fine-grained token for classification. As analyzed in (Brüggemann et al., 2021), this quantization only changes the way of predicting regression task, but does not contribute to the learning performance. For depth estimation, we follow the setting in (Brüggemann et al., 2021; Li et al., 2018), and divide the range of depth values into several logarithmic bins. Our predicted task logits score $\boldsymbol{G_{i}^{\prime}}$ is used to calculate the soft-weighted sum with each bin and produce final task predictions accordingly. For surface normal estimation, we follow (Brüggemann et al., 2021; Zeisl et al., 2014) and use K-means to learn several unit normal vectors, which serve as clustering centers, and they are also used to generate predictions with $\boldsymbol{G_{i}^{\prime}}$. This process can be expressed as:
(7) $\boldsymbol{\hat{Y_{i}}}=\operatorname{Sum}\left(\boldsymbol{c_{i}}^{\top}\times\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$
where $\boldsymbol{c_{i}}\in\mathbb{R}^{C_{p}}$ represents the center of each bin. In our experiments, the numbers of depth bins and normal cluster centers on NYUD-v2 are 303030 and 202020, respectively. For Cityscapes, we consider 100100100 depth bins as the Cityscapes dataset is captured from outdoor scenarios and have more significant changes in depth. For PASCAL-Context, we select 404040 different unit verctors uniformly from the space to serve as normal cluster centers.
(8) $\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right),$
(9) $\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}},{\boldsymbol{Y_{i}^{oh}}}\right)=-\boldsymbol{Y_{i}^{oh}}\cdot\operatorname{log}\left(\operatorname{Softmax}(\boldsymbol{G_{i}^{\prime}})\right),$
and the overall mixture loss for each task can be written as:
(10) $\operatorname{L_{i}}(\cdot)=\lambda_{i}\operatorname{L_{i}^{reg}}\left(\boldsymbol{\hat{Y_{i}}},{\boldsymbol{Y_{i}}}\right)+\operatorname{L_{i}^{cls}}\left(\boldsymbol{G_{i}^{\prime}},{\boldsymbol{Y_{i}^{oh}}}\right),$
A.4. Multi-scale Global Task Token Learning
For the implementations on NYUD-v2 and Cityscapes, we use the SegNet (Badrinarayanan et al., 2017) to serve as the shared backbone, which produces single-scale shared features for per-task decoding. However, on PASCAL-Context, we use ResNet-18 (He et al., 2016) as the shared backbone, which can produce multi-scale shared features for decoding. Directly fusing them ignores the various task information maintained in different scales. Thus, we propose to learn the global task tokens on different scales in order to learn more comprehensive task relations.
(11) $\boldsymbol{\theta_{i}}^{(j)}=\boldsymbol{W^{s\to m}_{i}}^{(j)}\times\boldsymbol{\theta_{i}},\quad j=0,1,2,3,$
[…]
(13) $\boldsymbol{\Theta}^{(j)}$ $=\left[\boldsymbol{\theta_{1}}^{(j)};\boldsymbol{\theta_{2}}^{(j)};\cdots;\boldsymbol{\theta_{T}}^{(j)}\right]^{\top}.$
[…]
(14) $\boldsymbol{F_{i}^{\prime}}$ $=\operatorname{Conv_{i}^{1\times 1}}\left(\left[\boldsymbol{F_{i}^{\prime}}^{(0)};\boldsymbol{F_{i}^{\prime}}^{(1)};\boldsymbol{F_{i}^{\prime}}^{(2)};\boldsymbol{F_{i}^{\prime}}^{(3)}\right]\right),$
(15) $\boldsymbol{\theta_{i}^{\prime}}$ $=\sum_{j=0}^{3}\left(\boldsymbol{W^{m\to s}_{i}}^{(j)}\times\boldsymbol{\theta_{i}^{\prime}}^{(j)}\right).$
Figure 9. Illustrations of Single-scale Global Task Token Learning (a) and Multi-scale Global Task Token Learning (b). For the multi-scale features produced by the shared backbone, we use linear layers to produce corresponding task tokens for each scale. Then, in each scale, we query task features and conduct inter-task attention to gain multi-task multi-scale representations. The multi-scale features and tokens are fused before intra-task attention to transfer cross-scale information.
A.5. Broader Applicability and Future Directions
The hierarchical design of our HiTTs is highly generalizable and not limited to the currently studied tasks. Its principles can be readily extended to other fundamental computer vision tasks, such as object detection (Bochkovskiy et al., 2020; Zhu et al., 2020; Cai et al., 2023; Ren et al., 2024; Wang et al., 2022a) and instance segmentation (Chen et al., 2023; Cheng et al., 2022; Schult et al., 2022; Gu et al., 2024), which fundamentally rely on rich, multi-scale pixel representations. The experiment on the PASCAL-Context dataset partially demonstrates this potential, where HiTTs effectively learns from both coarse-grained semantic labels and fine-grained human part annotations simultaneously, showcasing its robust joint-learning capabilities across different granularities.
Furthermore, the core concepts of HiTTs extend beyond the visual domain. Its ability to build cross-task relations upon multi-level representations makes it a promising approach for tasks like hierarchical text classification (Sun and Lim, 2001; Zhou et al., 2020; Stein et al., 2019; Wang et al., 2024, 2022b) or 3D scene-level parsing (Peng et al., 2023; Wang et al., 2025a; Jaritz et al., 2019). Moreover, the unified token-feature interaction mechanism, based on the Transformer architecture, is inherently compatible with diverse data types, paving the way for future exploration in multi-modal learning environments (Lin et al., 2023; Liu et al., 2023; Bai et al., 2023; Wang et al., 2023, 2025b). This adaptability underscores the broad potential of our hierarchical token-based approach for complex multi-task and multi-modal problems.
Appendix B Implementation Details
B.1. Dataset
@@
PASCAL-Context. PASCAL-Context (Everingham et al., 2009) contains 4998 and 5105 images for training and testing respectively, which also have pixel-level annotations for semantic segmentation, human-parts segmentation and semantic edge detection. Additionally, we also consider surface normal estimation and saliency detection distilled by (Maninis et al., 2019). We use Adam optimizer with learning rate $2\times 10^{-5}$, and weight decay $1\times 10^{-6}$, and train for 100100100 epochs with batch size 666. We update the learning rate with polynomial strategy and $\gamma=0.9$ for the power factor.
NYUD-v2. NYUD-v2 (Silberman et al., 2012) contains 795 and 654 RGB-D indoor scene images for training and testing respectively. We use the 13-class semantic annotations which is defined in (Couprie et al., 2013), the truth depth annotations recorded by Microsoft Kinect depth camera, and surface normal annotations which are produced in (Eigen and Fergus, 2015). Following the setting in (Liu et al., 2019; Li et al., 2022), we use 288×384288$\times$ 384288 × 384 image resolution to speed up training. We use Adam optimizer with a learning rate of $1\times 10^{-4}$, and train for 400 epochs with batch size 888. We update the learning rate every 100100100 epoch with $\gamma=0.5$ as the multiplying factor.
Cityscapes. Cityscapes (Cordts et al., 2016) contains 2975 and 500 street-view images for training and testing respectively. We used the projected 7-class semantic annotations from (Liu et al., 2019), and disparity maps to serve as depth annotations. Following the setting in (Liu et al., 2019; Li et al., 2022), we use 128×256128$\times$ 256128 × 256 image resolution to speed up training. The optimizer and learning rate scheduler are set as the same as NYUD-v2.
B.2. Data Preprocessing
We follow the setting of (Li et al., 2022; Liu et al., 2019; Vandenhende et al., 2021; Ye and Xu, 2024) to process training data. For NYUD-v2 and Cityscapes, we use random scaling, cropping and horizon flipping for data augmentation following (Li et al., 2022; Liu et al., 2019). For PASCAL-Context, we follow (Ye and Xu, 2024) and use random scaling, cropping, horizon flipping and photometric distortion for data augmentation. Apart from the two partial settings (one-label and random-labels) we mentioned in Sec.4.1, we also consider two extra settings: (iii) full-labels: each image has labels on every task; (iv) few-shot: one task has only very few labels while other tasks are fully supervised. Both of the extra settings will furthermore show the effectiveness of our method.
B.3. Model Setting
We use SegNet (Badrinarayanan et al., 2017) as the image backbone for our experiments on NYUD-v2 and Cityscapes; and we use ResNet-18 as the image backbone, Atrous Spatial Pyramid Pooling (ASPP) (Chen et al., 2018b) as the task-specific decoding heads. For the threshold $\tau_{i}$ which selects binary mask $\boldsymbol{M_{i}^{p}}$, we use {Semseg: 0.9, Depth: 0.45, Normal: 0.6 } for NYUD-v2 one-label setting, {Semseg: 0.9, Depth: 0.7, Normal: 0.7} for NYUD-v2 random-labels setting. {Semseg: 0.9, Depth: 0.5} for Cityscapes one-label setting, and {Semseg: 0.9, Parsing: 0.85, Normal: 0.7, Sal: 0.7, Edge: 0.9} for PASCAL-Context one-label and random-labels settings.